X Company has two production departments, A and B. The following is budgeted information for all of its products in 2019, and actual information for one of them, Product X:

Department A All Products Product X
   Overhead $5,800,000 --   
   Direct labor hours 50,000 825  
   Machine hours 100,000 1,070  
   Units produced 59,000 810  
Department B
   Overhead $1,900,000 --   
   Direct labor hours 40,000 159  
   Machine hours 120,000 820  
   Units produced 44,000 810  

Using a single-driver allocation system, with machine hours as the driver, how much overhead was allocated to Product X [round overhead rate(s) to two decimal places]?
Answer #1

Total estimated overheads = Estimated overheads of Department A + Estimated overheads of Department B

= 5,800,000+1,900,000

= $7,700,000

Total estimated machine hours = Estimated machine hours of Department A + Estimated machine hours of Department B

= 100,000+120,000

= $220,000

Predetermined overhead rate = Total estimated overheads / Total estimated machine hours

= 7,700,000/220,000

= $35

Machine hours used by Product X = Machine hours use in Department A + Machine hours use in Department B

= 1,070+820

= 1,890

Overhead allocated to product X = Machine hours used by Product X x Predetermined overhead rate

= 1,890 x 35

= $66,150

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks